OpenStudy (anonymous):

Paytrust.com is an online bill payment service that charges $5 plus o.50 cents per bill. Thomas used this service for 6 months and paid a total of $48.50. How many bills were paid during that period?

OpenStudy (texaschic101):

5 + .50b = 48.50 .50b = 48.50 - 5 .50b = 43.50 b = 43.50/.50 b = 87 check... 5 + .50(87) = 48.50 5 + 43.50 = 48.50 48.50 = 48.50 (correct) If I did this correctly, Thomas paid 87 bills. You might need to get a second opinion on this.

OpenStudy (texaschic101):

As long as the service doesn't charge $5 per month, which it doesn't state that, then I think it is correct.

OpenStudy (anonymous):

thank you, I thought i had it right but wasnt sure.

OpenStudy (texaschic101):

no problem...glad to help :)